Steve Conyers and Chelsy Bernard formed a partnership, dividing income as follows: Annual salary allowance to Bernard of $92,340. Interest of 5% on each partner's capital balance on January 1. Any remaining net income divided to Conyers and Bernard, 1:2. Conyers and Bernard had $81,000 and $75,000, respectively, in their January 1 capital balances. Net income for the year was $162,000. How much is distributed to Conyers and Bernard? Note: Compute partnership share.

Conyers: $

Bernard: $
Answer #1

Total amount of salary and interest payable to both the partners

= 92,340 + (5%*81,000) + (5%*75,000)

= 100,140

Remaining balance = 162,000 - 100,140 = 61,860

So, partnership share will be

Conyers = (5%*81,000) + (61,860*1/3) = 24,670

Bernard = (5%*75,000) + (61,860*2/3) + 92,340 = 137,330